# LLMs can Find Mathematical Reasoning Mistakes by Pedagogical Chain-of-Thought

### v1

@@ 2.3 Detecting Reasoning Errors with LLMs
Figure 3: The proposed PedCoT prompting method which evolved from left two columns of pedagogical ability levels and principles for prompt design. We design the instruction templates for PedCoT prompts that coordinate with the two-stage interaction process (TIP). The bold parts are keywords to embody the corresponding educational meanings. We refine the wording of PedCoT prompts to avoid misunderstanding by LLMs through numerous empirical attempts. The contents in [] are variables.
@@ 3.1 Problem Definition
Given a question (i.e., math problem) $Q$ and a trace of the top $i$ answer steps $A_{i}=\{a_{1},...,a_{i}\}$, as variables, they are integrated into predefined instruction templates $T=\{T_{j}^{k}\}$ and then the combined prompts can be represented by $P=\{P_{j}\}$, where $j$ means the $j$-th stage of request on a pre-selected LLM (e.g., GPT-4) and $k$ corresponds to the $k$-th pedagogical principle. In this work, we set the $J=2$ and $K=3$ since we only need two stages of requests on LLM to complete the mistake finding task and we leverage three pedagogical principles for the time being. Therefore, at Stage-1, the input prompt $P_{1}$ is combined with $Q$, $A_{i-1}$ and $T_{1}$ to request on the LLM to obtain the intermediate content $G^{(k)}$. The process is named as Regenerate and is introduced in the following section. At Stage-2, the input prompt $P_{2}$ is combined with $Q$, $A_{i}$, $T_{2}$ and $G$ to request on the LLM again to extract some content $E^{(k)}$ and compare the consistency between the regenerated content and extracted content. This process is called as Extract-Compare. The LLM can finally output a principle mistake label in terms of principles which is defined as:
@@
where correct-and-aligned means the extracted contents $E^{(k)}$ are correct and aligned with the regenerated contents $G^{(k)}$. reasonable-but-incomplete means the extracted contents $E^{(k)}$ are partially correct but not completely aligned with the regenerated contents $G^{(k)}$. contradiction-found means the extracted contents $E^{(k)}$ are totally contrary to the regenerated contents $G^{(k)}$. nothing-extracted means no related content is extracted from the current answer step.
@@
$L_{step}^{(i)}=\begin{cases}1,&\text{if}\forall k\in\{1,2,3\},st.L_{principle}^{(k)}\neq\text{contradiction-found},\\ 0,&\text{if}\exists k\in\{1,2,3\},st.L_{principle}^{(k)}=\text{contradiction-found}.\end{cases}$
@@
$L_{trace}=\begin{cases}1,&\text{if}\forall i\in\{1,...,n\},st.L_{step}^{(i)}=1,\\ 0,&\text{if}\exists i\in\{1,...,n\},st.L_{step}^{(i)}=0,\end{cases}$
@@ 3.2 Analysis
Pedagogical Principles for Prompt Design
@@ 4.1 Two-stage Interaction Process
Stage-2: Extract-Compare
@@
BIG-Bench Mistake PRM800K
# Total Traces 300 300
# Total Steps 1,506 3,736
# Avg. Trace Length 5.02 12.45
# Avg. Steps of 1st Mistake 2.44 4.89
# Correct Traces(%) 62(20.67%) 85(28.33%)
Table 1: Statistics of testing datasets.
@@ 4.2 Pedagogical Chain-of-Thought Prompts
Based on pedagogical principles mentioned in Section 3.2, we develop the Pedagogical Chain-of-Thought (PedCoT) prompts for each stage of TIP. Each prompt consists of the instruction templates (i.e., $T_{1}$ or $T_{2}$) and the variables (i.e., $Q$, $A$ and/or $G$). As shown in Figure 3, the right two columns list the grounded contents of developed prompts for each stage. The bold parts are keywords to embody the corresponding educational meanings. We refine the wording of grounded prompts to avoid misunderstanding by LLMs through numerous empirical attempts.
@@ 5.1 Datasets
To evaluate our proposed PedCoT, we collect two public datasets containing step-level correctness labels for mathematical problems with different difficulties. The statistics of the two datasets are summarized in Table 1.
@@
Tyen et al. (2023): The original dataset consists 2,186 reasoning traces generated by PALM-2 Anil et al. (2023), encompassing five reasoning tasks derived from diverse domains. With an emphasis on empowering LLMs to identify reasoning errors within the realm of mathematics, we chose samples exclusively related to this field. In total, these samples encompass 1,506 reasoning steps across 300 answer traces. The type of mathematical problem consistently involves multi-step arithmetic calculations. These are relatively straightforward to solve and necessitate basic mathematical concepts and problem-solving strategies, such as the fundamental rules of arithmetic. As shown in Table 1, some correct traces are involved to avoid the bias that LLMs may treat every trace as faulty.
LLMs CoT methods MF. Acc. Avg. F1 Cls. Acc. $P_{+}$ $R_{+}$ $F_{+}$ $P_{-}$ $R_{-}$ $F_{-}$
GPT-4 Direct-Step 42.67 78.67 - - - - - - -
@@
Table 2: Comparison of various CoT methods on the BIG-Bench Mistake dataset. The scores of Direct-Step are from the original paper.
PRM800K
Lightman et al. (2023): The original dataset contains about 800,000 step-level labels over 75,000 solutions for math word problems in the MATH dataset Hendrycks et al. (2021), which consists of 12,500 problems from high-school math competitions. The problems necessitate a comprehensive understanding of intricate mathematical concepts and problem-solving methodologies. We adopt this challenging dataset to verify the effectiveness and robustness of our method. We randomly select a set of 300 pairs of problems and their affiliated reasoning answer traces [^3] Accessible on github.com/HaoyuanPeng/PedCoT-IJCAI24/, containing 85 correct traces and 3,736 reasoning steps in total, shown in Table 1. The steps in the PRM800K dataset are categorized as positive, negative, and neutral. To use the proposed method, we adjust the labels to align with our label schema. Specifically, the positive or neutral steps are treated as right (i.e., $L_{step}=1$) while the negative steps are labeled as wrong (i.e., $L_{step}=0$) in our experiments.
5.2 Models, Baselines, and Metrics
Models
Baselines
Metrics
5.3 Results and Analysis
Table 2 and Table 3 report the comparison among different zero-shot CoT methods applied with GPT-4 and GPT-4 Turbo on the two datasets. Consistent with previous research, directly prompting the LLMs to output a single word for the correctness of each reasoning step (i.e., Direct-Step) is ineffective even with the state-of-the-art LLMs. Vanilla Two-stage Prompting allows LLMs to generate several analyses before being prompted again for the final label word, leading to a slightly better performance than Direct-Step. Zero-shot CoT and Plan-and-Solve Prompting explicitly instruct LLMs to generate intermediate analyses, and show improvements compared to Vanilla Two-stage Prompting. The above results indicate that the more analysis contents before giving the final judgment are generated by prompting LLMs, the better performance is obtained, especially with GPT-4 Turbo.
LLMs CoT methods MF. Acc. Avg. F1 Cls. Acc. $P_{+}$ $R_{+}$ $F_{+}$ $P_{-}$ $R_{-}$ $F_{-}$
GPT-4 Direct-Step 33.33 67.10 66.00 42.20 54.12 47.42 79.58 70.70 74.88
@@ 5.3 Results and Analysis
Table 3: Comparison of various CoT methods on the PRM800K dataset.
@@ 5.4 Ablation Study
Ablating prompt contents of Principle 1 (Remember)
@@
Ablating prompt contents of Principle 2 (Understand)
@@
Ablating prompt contents of Principle 3 (Apply)
@@
Table 5: Two-stage PedCoT vs. one-stage variant on GPT-4.

### v2

@@ 2.3 Detecting Reasoning Errors with LLMs
Table 1: The proposed PedCoT prompting method which evolved from left two columns of pedagogical ability levels and principles for prompt design. We design the instruction templates for PedCoT prompts that coordinate with the two-stage interaction process (TIP). The bold parts are keywords to embody the corresponding educational meanings. We refine the wording of PedCoT prompts to avoid misunderstanding by LLMs through numerous empirical attempts. The contents in [] are variables.
@@ 3.1 Problem Definition
$L_{step}^{(i)}=\begin{cases}1,&\text{if}\quad\forall k\in\{1,2,3\},st.L_{principle}^{(k)}\neq\text{contradiction-found},\\ 0,&\text{if}\quad\exists k\in\{1,2,3\},st.L_{principle}^{(k)}=\text{contradiction-found}.\end{cases}$
@@
$L_{trace}=\begin{cases}1,&\text{if}\quad\forall i\in\{1,...,n\},st.L_{step}^{(i)}=1,\\ 0,&\text{if}\quad\exists i\in\{1,...,n\},st.L_{step}^{(i)}=0,\end{cases}$
@@ 3.2 Analysis
Pedagogical Principles for Prompt Design.
@@ 4.1 Two-Stage Interaction Process
Stage-2: Extract-Compare.
@@ 4.2 Pedagogical Chain-of-Thought Prompts
Based on pedagogical principles mentioned in Section 3.2, we develop the Pedagogical Chain-of-Thought (PedCoT) prompts for each stage of TIP. Each prompt consists of the instruction templates (i.e., $T_{1}$ or $T_{2}$) and the variables (i.e., $Q$, $A$ and/or $G$). As shown in Table 1, the right two columns list the grounded contents of developed prompts for each stage. The bold parts are keywords to embody the corresponding educational meanings. We refine the wording of grounded prompts to avoid misunderstanding by LLMs through numerous empirical attempts.
@@ 5.1 Datasets
To evaluate our proposed PedCoT, we collect two public datasets containing step-level correctness labels for mathematical problems with different difficulties.
@@
Tyen et al. (2023): The original dataset consists 2,186 reasoning traces generated by PALM-2 Anil et al. (2023), encompassing five reasoning tasks derived from diverse domains. With an emphasis on empowering LLMs to identify reasoning errors within the realm of mathematics, we chose samples exclusively related to this field. In total, these samples encompass 1,506 reasoning steps across 300 answer traces. The type of mathematical problem consistently involves multi-step arithmetic calculations. These are relatively straightforward to solve and necessitate basic mathematical concepts and problem-solving strategies, such as the fundamental rules of arithmetic. 62 correct traces are involved to avoid the bias that LLMs may treat every trace as faulty.
Datasets LLMs CoT methods MF. Acc. Avg. F1 Cls. Acc. $P_{+}$ $R_{+}$ $F_{+}$ $P_{-}$ $R_{-}$ $F_{-}$
BIG-Bench Mistake GPT-4 Direct-Step 42.67 78.67 - - - - - - -
@@
PRM800K GPT-4 Direct-Step 33.33 67.10 66.00 42.20 54.12 47.42 79.58 70.70 74.88
@@
Table 2: Comparison of various CoT methods on the BIG-Bench Mistake and PRM800K datasets. The scores are from the original paper.
PRM800K
Lightman et al. (2023): The original dataset contains about 800,000 step-level labels over 75,000 solutions for math word problems in the MATH dataset Hendrycks et al. (2021), which consists of 12,500 problems from high-school math competitions. The problems necessitate a comprehensive understanding of intricate mathematical concepts and problem-solving methodologies. We adopt this challenging dataset to verify the effectiveness and robustness of our method. We randomly select a set of 300 pairs of problems and their affiliated reasoning answer traces [^3] Accessible on github.com/HaoyuanPeng/PedCoT-IJCAI24/, containing 85 correct traces and 3,736 reasoning steps in total. The steps in the PRM800K dataset are categorized as positive, negative, and neutral. To use the proposed method, we adjust the labels to align with our label schema. Specifically, the positive or neutral steps are treated as right (i.e., $L_{step}=1$) while the negative steps are labeled as wrong (i.e., $L_{step}=0$) in our experiments.
5.2 Models, Baselines, and Metrics
Models.
Baselines.
Metrics.
5.3 Results and Analysis
Table 2 reports the comparison among different zero-shot CoT methods applied with GPT-4 and GPT-4 Turbo on the two datasets. Consistent with previous research, directly prompting the LLMs to output a single word for the correctness of each reasoning step (i.e., Direct-Step) is ineffective even with the state-of-the-art LLMs. Vanilla Two-stage Prompting allows LLMs to generate several analyses before being prompted again for the final label word, leading to a slightly better performance than Direct-Step. Zero-shot CoT and Plan-and-Solve Prompting explicitly instruct LLMs to generate intermediate analyses, and show improvements compared to Vanilla Two-stage Prompting. The above results indicate that the more analysis contents before giving the final judgment are generated by prompting LLMs, the better performance is obtained, especially with GPT-4 Turbo.
@@ 5.4 Ablation Study
Ablating prompt contents of Principle 1 (Remember).
@@
Ablating prompt contents of Principle 2 (Understand).
@@
Ablating prompt contents of Principle 3 (Apply).
@@
Table 4: Two-stage PedCoT vs. its one-stage variant on GPT-4.
@@ Acknowledgments
Appendix A Using PedCoT on GPT-3.5-Turbo
Table 5 shows the comparison between PedCoT prompting and Direct-Step prompting with GPT-3.5-Turbo. PedCoT outperforms Direct-Step prompting by a significant margin on both datasets. Particularly, on the BIG-Bench Mistake dataset, PedCoT improves the mistake finding accuracy from 26.00% to 75.67%, and the average F1 from 64.08% to 87.20%. On the PRM800K dataset, PedCoT achieved improvements of 16.67%, 13.20%, and 13.33% in MF. Acc., Avg. F1, and Cls. Acc. respectively. This experiment can indicate that our PedCoT method is also effective with an earlier low-version LLM.
Datasets CoT methods MF. Acc. Avg. F1 Cls. Acc. $P_{+}$ $R_{+}$ $F_{+}$ $P_{-}$ $R_{-}$ $F_{-}$
BIG-Bench Mistake Direct-Step 26.00 64.08 - - - - - - -
PedCoT (ours) 75.67 87.20 88.00 79.55 56.45 66.04 89.45 96.22 92.71
PRM800K Direct-Step 20.00 63.04 62.67 35.16 37.65 36.36 74.64 72.56 73.58
PedCoT (ours) 36.67 76.24 76.00 57.14 61.18 59.09 84.21 81.86 83.02
Table 5: Comparison of our PedCoT with baseline when using GPT-3.5-Turbo.
Appendix B The Complete PedCoT Prompts and LLM Responses of the Example in Main Body’s Figure 2
We present a complete two-stage interaction process (TIP) of PedCoT strategy with a LLM (e.g., GPT-4 Turbo) in Table LABEL:appendix:example. The background colors are consistent with main body’s Figure 2, that is, the texts with a grey background represent the instruction template $T_{j}$, the texts with a purple background represent the question $Q$, the texts with a green background represent the answer steps $\{a_{i}\}$, the texts with a yellow background represent the outputs $\{G^{(k)}\}$ from the LLM at Stage-1, the texts with a blue background represent the outputs $\{E^{(k)}\}$ extracted by the LLM at Stage-2, and the texts with a red background represent the comparison and principle mistake labels $L_{principle}^{(k)}$, where $j$ means the $j$-th stage, $i$ means the $i$-th answer step, and $k$ corresponds to the $k$-th pedagogical principle. The uncolored texts are just notes to segment the input and output of two stages and are not part of either the input prompts or LLM responses.
[…]
Table 6: An example of the complete prompt texts and LLM responses, corresponding to the main body’s Figure 2.
Stage-1 Prompt: $P_{1}=\{Q,A_{i-1},T_{1}\}$
You are given a math problem and several initial steps of the solution process.
Question: Let $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}.$ Find the range of $f(x).$ All functions are in radians.
Initial steps:
(step 1) I notice that the domain of both $\arccos$ and $\arcsin$ is the interval $[-1,1]$, so I can start by assuming that $x$ is in that interval.
Execute the following instructions sequentially. 1. List the mathematical concepts that should be applied in the next step for solving the problem. 2. List the key analyses that should be derived in the next step for solving the problem. 3. List the mathematical expressions that have not been fully computed yet in the question and the initial steps. Then carefully calculate their corresponding results step by step.
Stage-1 Response: $G=\{G^{(1)},G^{(2)},G^{(3)}\}$
1. Mathematical Concepts to Apply:
- Trigonometric identities, specifically the Pythagorean identity: $\sin^{2}(\theta)+\cos^{2}(\theta)=1$.
- The fact that $\arcsin(x)$ and $\arccos(x)$ are inverse functions of $\sin(\theta)$ and $\cos(\theta)$, respectively, and that for a given value of $x$, $\sin(\arcsin(x))=x$ and $\cos(\arccos(x))=x$.
- The properties of the $\arcsin$ and $\arccos$ functions, including their ranges: $\arcsin(x)\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\arccos(x)\in[0,\pi]$.
- Calculus concepts may be used to find the extrema of the function $f(x)$ if necessary.
2. Key Analyses for the Next Step:
- Recognize the relationship between the functions $\arccos(x)$ and $\arcsin(x)$.
- Use the Pythagorean identity for sine and cosine to relate $\arccos(x)$ and $\arcsin(x)$.
- Determine the maximum and minimum values of the function $f(x)$ within the domain of $x\in[-1,1]$.
3. Mathematical Expressions to Compute:
- We need to compute the expression for $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}$ using the relationship between $\arccos(x)$ and $\arcsin(x)$.
Let’s compute the expression step by step:
Since $\arccos(x)$ and $\arcsin(x)$ are related by the identity $\arccos(x)+\arcsin(x)=\frac{\pi}{2}$ for $x\in[-1,1]$, we can use this to simplify $f(x)$.
We have: $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}$ $f(x)=\left(\frac{\pi}{2}-\arcsin(x)\right)^{2}+(\arcsin x)^{2}$ $f(x)=\frac{\pi^{2}}{4}-\pi\arcsin(x)+(\arcsin(x))^{2}+(\arcsin x)^{2}$ $f(x)=\frac{\pi^{2}}{4}-\pi\arcsin(x)+2(\arcsin(x))^{2}$
Stage-2 Prompt: $P_{2}=\{Q,A_{i},G,T_{2}\}$
You are given a math problem and several initial steps of the solution process.
Question: Let $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}.$ Find the range of $f(x).$ All functions are in radians.
Initial steps:
(step 1) I notice that the domain of both arccos and arcsin is the interval [-1, 1], so I can start by assuming that x is in that interval.
The actual next step is:
(step 2) Next, I recall that arccos and arcsin are inverse functions, meaning that arccos(arcsin(x)) = x and arcsin(arccos(x)) = x for any x in [-1, 1].
Execute the following instructions sequentially. 1. First extract the mathematical concepts employed in the actual next step. Below is the correct mathematical concepts should be applied in the next step. Give a detailed comparison between the extracted concepts and the correct mathematical concepts.
Mathematical Concepts to Apply:
- Trigonometric identities, specifically the Pythagorean identity: $\sin^{2}(\theta)+\cos^{2}(\theta)=1$.
- The fact that $\arcsin(x)$ and $\arccos(x)$ are inverse functions of $\sin(\theta)$ and $\cos(\theta)$, respectively, and that for a given value of $x$, $\sin(\arcsin(x))=x$ and $\cos(\arccos(x))=x$.
- The properties of the $\arcsin$ and $\arccos$ functions, including their ranges: $\arcsin(x)\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\arccos(x)\in[0,\pi]$.
- Calculus concepts may be used to find the extrema of the function $f(x)$ if necessary.
Finally, output a label from the options (correct-and-aligned, reasonable-but-incomplete, nothing-extracted, and contradiction-found) to categorize the use of mathematical concepts for the actual next step.
2. First extract the key analyses present in the actual next step. Below is the correct analyses should be derived in the next step. Give a detailed comparison between the actual analyses and the correct analyses.
Key Analyses for the Next Step:
- Recognize the relationship between the functions $\arccos(x)$ and $\arcsin(x)$.
- Use the Pythagorean identity for sine and cosine to relate $\arccos(x)$ and $\arcsin(x)$.
- Determine the maximum and minimum values of the function $f(x)$ within the domain of $x\in[-1,1]$.
Finally, output a label from the options (correct-and-aligned, reasonable-but-incomplete, nothing-extracted, and contradiction-found) to categorize the problem solving approach for the actual next step.
3. First extract the calculation results present in the actual next step. Below is the correct calculation to be done in the next step. If any calculation results are extracted from the actual next step, list their corresponding results in the correct calculations, and give a detailed comparison between the extracted calculation results and correct results. Mathematical Expressions to Compute:
- We need to compute the expression for $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}$ using the relationship between $\arccos(x)$ and $\arcsin(x)$.
Let’s compute the expression step by step:
Since $\arccos(x)$ and $\arcsin(x)$ are related by the identity $\arccos(x)+\arcsin(x)=\frac{\pi}{2}$ for $x\in[-1,1]$, we can use this to simplify $f(x)$.
We have: $f(x)=(\arccos x)^{2}+(\arcsin x)^{2}$ $f(x)=\left(\frac{\pi}{2}-\arcsin(x)\right)^{2}+(\arcsin x)^{2}$ $f(x)=\frac{\pi^{2}}{4}-\pi\arcsin(x)+(\arcsin(x))^{2}+(\arcsin x)^{2}$ $f(x)=\frac{\pi^{2}}{4}-\pi\arcsin(x)+2(\arcsin(x))^{2}$
Finally, output a label from the options (correct-and-aligned, reasonable-but-incomplete, nothing-extracted, and contradiction-found) to categorize the correctness of calculations for the actual next step.
Stage-2 Response: $E=\{E^{(1)},E^{(2)},E^{(3)}\}$ and $L_{principle}^{(k)}$
1. Mathematical Concepts to Apply:
The actual next step does not employ the correct mathematical concepts. The correct concepts involve using the Pythagorean identity and understanding the properties of the arcsin and arccos functions, including their ranges. The actual next step does not mention the Pythagorean identity or the ranges of the functions. Therefore, the label for the use of mathematical concepts is ”contradiction-found.”
2. Key Analyses for the Next Step:
The actual next step incorrectly states that arccos and arcsin are inverse functions in a way that suggests $arccos(arcsin(x))=x$ and $arcsin(arccos(x))=x$ for any $x$ in $[-1,1]$, which is not true. The correct relationship is that for a given angle $\theta$, $sin(arccos(\theta))$ and $cos(arcsin(\theta))$ will give the corresponding leg lengths of a right triangle where $\theta$ is one of the non-right angles. The correct analyses should recognize the relationship between $arccos(x)$ and $arcsin(x)$ and use the Pythagorean identity to relate them. The actual next step does not do this, and therefore, the label for the problem-solving approach is ”contradiction-found.”
3. Mathematical Expressions to Compute:
No calculations are extracted from the actual next step. The correct calculation involves using the identity $arccos(x)+arcsin(x)=\pi/2$ and then finding the range of $f(x)$ by considering the maximum and minimum values of $arcsin(x)$ within its domain. Since the actual next step does not provide any calculation results, the label for the correctness of calculations is ”nothing-extracted.”